The measure of one of two complementary angles is three more than twice the other. The smaller equals _ ?
Answer by Earlsdon(6294)

Let the two angles be A and B.
A+B = 90 "...two complementary angles..."
A = 2B+3 "The measure of one of two complementary angles is three more than twice the other." Substitute this for A in the first equation and solve for angle B.
(2B+3)+B = 90 Simplify.
3B+3 = 90 Subtract 3 from both sides.
3B = 87 Divide both sides by 3.
B = 29 degrees.
A = 2(29)+3
A = 61 degrees.
The smaller angle equals 29 degress.